OpenStudy (anonymous):

Simplify the following expression. 2x - 2y + 5z - 2x - y + 3z

OpenStudy (anonymous):

Combine like terms. Remember to keep the sign in front of each term. 2x - 2y + 5z - 2x - y + 3z 2x - 2x = 0x -2y - y = -3y 5z + 3z = 8z So your equation is -3y + 8z